OpenStudy (anonymous):

Andrea's fish tank holds 30 gallons of water. Andrea estimates that every inch of a fish's length requires a gallon of water. She wants to purchase some corydoras that are 2 inches long, angelfish that are 5 inches long, and fancy guppies that are 3 inches long. Write an equation for the situation.

OpenStudy (anonymous):

A) 2x + 5y + 3z = 30 B) x + y + z = 30 C) 2x + 5y + 3z = 300 D) x + y + z = 100

OpenStudy (tyteen4a03):

A would be the answer, assuming that you don't mean "extra water". (But what kind of fish would need a gallon of water anyway?)

OpenStudy (anonymous):

lol i don't know it was a homework question

OpenStudy (anonymous):

thanks